OpenStudy (anonymous):

M(9, 8) is the midpoint of mc136-1.jpg. The coordinates of S are (10, 10). What are the coordinates of R

OpenStudy (anonymous):

to find mid point formula is.. M(x,y)= (x1+x2/2 ,y1+y2/2)

OpenStudy (anonymous):

so it will be.. (9,8)=(10+x2/2 ,10+y2/2) compare both saides coordinates. 9=10+x2/2 =>10+x2=18 =>x2=18-10 =>x2=8 and 8=10+y2/2 =>10+y2=16 =>y2=16-10 =>y2=6 so points of R=(8,6)